10. Find the critical value tc for the level of confidence c and sample size n....

10.

Find the critical value tc for the level of confidence c and sample size n.

c = 0.90, n = 4

tc = ± 2.353

tc = ± 5.841

tc = ± 2.132

tc = ± 1.645
Answer #1

solution

n=4

df =n-1=4-1=3

= 1 - C = 1 - 0.90 = 0.1

/ 2 = 0.1 / 2 = 0.05

t /2,df = t0.05,3 = ± 2.353 using student t table